OpenStudy (anonymous):

Given the sequence –3, 0, 3, which of the following is equivalent to:

OpenStudy (anonymous):

\[\sum_{n=5}^{8}A -subscript- N\]

OpenStudy (anonymous):

54 102 144 162

OpenStudy (anonymous):

@primeralph @terenzreignz

OpenStudy (anonymous):

@Zarkon

OpenStudy (anonymous):

First term is -3, common difference is 0-(-3) = 3 So, the sequence is \[a_n = -3+3(n-1)\]

OpenStudy (anonymous):

so, we can do that plugging n-2, n-3 etc. into the next equations?

OpenStudy (anonymous):

[-3+3(n-1)]+[-3+3(n-2)] etc.?

OpenStudy (anonymous):

You need to find sum from the fifth term to the eighth term, so, find a5, a6, a7, a8 and add them together.

OpenStudy (anonymous):

n is the number of term

OpenStudy (anonymous):

that's right we start on 5 right? so like this?

OpenStudy (anonymous):

e.g. \[a_5 = -3+3(5-1)=...\]

OpenStudy (anonymous):

Start from 5, yes.

OpenStudy (anonymous):

[-3+3(n-5)]+[-3+3(n-6)]+[-3+3(n-7)]+[-3+3(n-8)]?

OpenStudy (anonymous):

(-3+3n-15)+(-3+3n-18)+ (-3+3n-21)+ (-3+3n-24)?

OpenStudy (anonymous):

am I doing this right?

OpenStudy (anonymous):

(3n-18)+(3n-21)+(3n-24)+(3n-x)+(3n-27)

OpenStudy (anonymous):

i'm not sure what to do from there

OpenStudy (anonymous):

No

OpenStudy (anonymous):

\[a_5 = -3+3(5-1)\]\[a_6 = -3+3(6-1)\]\[a_7 = -3+3(7-1)\]... Then, find \(a_5+a_6+a_7+a_8\)

OpenStudy (anonymous):

=−3+3(5−1) =−3+3(6−1) =−3+3(7−1) =-3+3(8-1)

OpenStudy (anonymous):

=-6+15 =-6+18 =-6+21 =-6+24 =9 =12 =15 =18 9+12+15+18

OpenStudy (anonymous):

54

OpenStudy (anonymous):

thank you so much guys